### v1

# Do massive neutrino states really exist?

Abstract
In neutrino physics, it is conventional to work with massive states corresponding to defined neutrino masses. However, in weak interaction processes, only flavor states participate, and the propagation can be described by an off-diagonal mass matrix. In this work, we operate directly with the non-diagonal mass matrix and employ a quantum field theory approach to describe neutrino oscillations within the plane-wave approximation, circumventing the use of wave packets. Within this framework, numerical estimates for the amplitude of a local process involving lepton interactions via virtual neutrino exchange yield values of order $10^{-53}$. For macroscopic processes with neutrino exchange, application of the matrix propagator reproduces the standard quantum mechanical neutrino oscillation probability formula while preserving relativistic invariance.
— …I didn’t ask you whether you believe that ghosts are seen, but whether you believe that they exist.
F. Dostoyevsky
<<Crime and Punishment>>
Translated by R. Pevear
and L. Volokhovsky
Introduction
Typically, the neutrino which is propagating for the long distances, is considered as a real particle and thus must exist in a mass eigenstate. This is achieved through the diagonalization of the mass matrix entering the respective Lagrangian. However, within the modern framework, it is treated as a virtual particle. This then enables the computation of these propagators without having to diagonalize the mass matrix.
In this work, we extend the QFT approach to processes with lepton number violation, analyzing a two-flavor model of Dirac neutrinos. Unlike the work of Kovalenko and $\check{\text{S}}$imkovic [3] , who derived diagonal propagators for massive neutrinos, we consider flavor-neutrino propagation which are described by a non-diagonal matrix propagator. A similar methodology was employed by M. Dvornikov [4] in analyzing oscillations in matter.
Next, we consider the mass matrix in the Chau-Keung parameterization of the PMNS matrix via Euler angles. We discuss how parameterization can affect the normal and inverted neutrino mass hierarchies.
We demonstrate that local weak processes (e.g., $l^{-}_{\alpha}l^{+}_{\beta}\to W^{+}W^{-}$) exhibit oscillation-sensitive contributions suppressed by $\mathcal{O}(10^{-79})$. For macroscopic diagrams with two-flavor transitions, the derived amplitude reproduces the canonical oscillation dependence
$P_{osc}\propto\sin^{2}\left(\cfrac{\Delta m^{2}L}{4E_{\text{avg}}}\right),$
where the oscillation phase is directly linked to neutrino masses. This formalism bridges QFT foundations and experimental interpretation while simultaneously confirming consistency with quantum mechanical results.
1 Dirac Neutrino Propagator
The Lagrangian for Dirac flavor neutrinos in a diagonalize form is given by:
$\mathcal{L}_{\nu}(x)=\bar{\nu}_{\alpha}(x)(i\hat{\partial}\cdot\delta_{\alpha\beta}-M_{\alpha\beta})\nu_{\beta}(x),$ (1.1)
where summation is over flavor states. The process of obtaining the propagator consists of inverting the matrix operator
$\mathbb{K}_{\alpha\beta}=i\hat{\partial}\cdot\delta_{\alpha\beta}-M_{\alpha\beta}.$ (1.2)
The propagator in momentum space, analogous to [1], is defined as $\mathbb{K}^{-1}(p)\equiv\mathbb{D}(p)$ :
$\mathbb{D}(p)=\frac{1}{\hat{p}-\mathbb{M}}.$ (1.3)
1.1 Two-Flavor Propagator
We consider the specific case of interaction between two flavors, electron ($e$) and muon ($\mu$), with the mass matrix:
$\mathbb{M}=\begin{pmatrix}m_{e}&m_{e\mu}\\ m_{e\mu}&m_{\mu}\end{pmatrix},$ (1.4)
having eigenvalues $m_{k}$.
The propagator in momentum space takes the form:
$\mathbb{D}(p)=\frac{(\hat{p}+m_{e})(\hat{p}+m_{\mu})-m_{e\mu}^{2}}{(p^{2}-m_{1}^{2})(p^{2}-m_{2}^{2})}\begin{pmatrix}m_{e}&m_{e\mu}\\ m_{e\mu}&m_{\mu}\end{pmatrix}.$ (1.5)
$m_{1}$ and $m_{2}$ are the neutrino masses and are expressed in terms of the mass matrix invariants as
$m_{k}^{2}=\cfrac{\text{tr}^{2}(\mathbb{M})}{2}-\det(\mathbb{M})+(-1)^{k}\ \cfrac{\text{tr}(\mathbb{M})}{2}\sqrt{\text{tr}^{2}(\mathbb{M})-4\det(\mathbb{M})},\quad k=1,2.$ (1.6)
It is convenient to represent it as
$\mathbb{D}(p)=\frac{(\hat{p}+\mathbb{M})\cdot\widetilde{\mathbb{A}}(p^{2})}{(p^{2}-m_{1}^{2})(p^{2}-m_{2}^{2})},$ (1.7)
where matrix $\widetilde{\mathbb{A}}(p^{2})$ is
$\widetilde{\mathbb{A}}(p^{2})=\begin{pmatrix}p^{2}+\det(\mathbb{M})-\text{tr}(\mathbb{M})\cdot m_{\mu}&\text{tr}(\mathbb{M})\cdot m_{e\mu}\\ \text{tr}(\mathbb{M})\cdot m_{e\mu}&p^{2}+\det(\mathbb{M})-\text{tr}(\mathbb{M})\cdot m_{e}\\ \end{pmatrix}.$ (1.8)
If $m_{e\mu}$ tends to 0 then
$\widetilde{\mathbb{A}}(p^{2})=\begin{pmatrix}p^{2}-m_{\mu}^{2}&0\\ 0&p^{2}-m_{e}^{2}\\ \end{pmatrix}$ (1.9)
and propagator (1.7) transform to diagonal form
$\lim\limits_{m_{e\mu}\to 0}\mathbb{D}(p)=\text{diag}\left(\frac{\hat{p}+m_{e}}{p^{2}-m_{e}^{2}},\quad\frac{\hat{p}+m_{\mu}}{p^{2}-m_{\mu}^{2}}\right)$ (1.10)
1.2 Propagator and mixing angle
It is interesting to express the resulting propagator (1.7) in terms of mixing angle and neutrino masses.
$\mathbb{M}=U\ \text{diag}(m_{1},m_{2})\ U^{\dagger},$ (1.11)
where $m_{1},\ m_{2}$ neutrino mass, and unitary rotation matrix is
$U=\begin{pmatrix}\cos\theta&\sin\theta\\ -\sin\theta&\cos\theta\\ \end{pmatrix}.$ (1.12)
Expressing the components of the propagator through these quantities, we obtain
$\begin{pmatrix}m_{e}&m_{e\mu}\\ m_{e\mu}&m_{\mu}\\ \end{pmatrix}=\begin{pmatrix}m_{1}\cos^{2}\theta+m_{2}\sin^{2}\theta&(m_{2}-m_{1})\cfrac{\sin(2\theta)}{2}\\ (m_{2}-m_{1})\cfrac{\sin(2\theta)}{2}&m_{1}\sin^{2}\theta+m_{2}\cos^{2}\theta\end{pmatrix}$ (1.13)
and
$\widetilde{\mathbb{A}}(p^{2})=\begin{pmatrix}(p^{2}-m_{1}^{2})\sin^{2}\theta+(p^{2}-m_{2}^{2})\cos^{2}\theta&\cfrac{\Delta m^{2}}{2}\sin(2\theta)\\ \cfrac{\Delta m^{2}}{2}\sin(2\theta)&(p^{2}-m_{1}^{2})\cos^{2}\theta+(p^{2}-m_{2}^{2})\sin^{2}\theta\\ \end{pmatrix},$ (1.14)
where $\Delta m^{2}\equiv m_{2}^{2}-m_{1}^{2}$.
As a result of substituting (1.13) and (1.14) into the (1.7), we obtain the full form of the propagator in terms of angles and neutrino masses
$\mathbb{D}(p)=\begin{pmatrix}D_{1}(p)\cdot\cos^{2}\theta+D_{2}(p)\cdot\sin^{2}\theta&\left[D_{2}(p)-D_{1}(p)\right]\cdot\cfrac{\sin(2\theta)}{2}\\ \left[D_{2}(p)-D_{1}(p)\right]\cdot\cfrac{\sin(2\theta)}{2}&D_{1}(p)\cdot\sin^{2}\theta+D_{2}(p)\cdot\cos^{2}\theta\\ \end{pmatrix},$ (1.15)
and $D_{1},\ D_{2}$ correspond to mass neutrino propagator
$D_{1}(p)=\cfrac{\hat{p}+m_{1}}{p^{2}-m_{1}^{2}},\quad\quad D_{2}(p)=\cfrac{\hat{p}+m_{2}}{p^{2}-m_{2}^{2}}.$ (1.16)
2 Three flavor mass matrix
Analyzing the propagator for three neutrino flavors in its off-diagonal form presents several complications. However, it is useful to trace the connection between the mass matrix and the mixing angles for three flavors. When mass matrix have only real elements, we have 6 independent variables that can be expressed through 6 parameters: 3 masses neutrino and 3 mixing angles. But we have a Hermitian mass matrix with nine different variables and (2.2) is expressed through the only seven independent parameters: 3 masses, 3 mixing angles and CP-phase $\delta$. We expect there are two relation between mass matrix elements. In the general case, the mass matrix is Hermitian:
$\mathbb{M}_{\text{flavor}}=\begin{pmatrix}m_{e}&m_{e\mu}&m_{e\tau}\\ m_{\mu e}&m_{\mu}&m_{\mu\tau}\\ m_{\tau e}&m_{\tau\mu}&m_{\tau}\end{pmatrix}.$ (2.1)
To represent it in the form
$\mathbb{M}_{\text{flavor}}=V\cdot\text{diag}(m_{1},m_{2},m_{3})\cdot V^{\dagger},$ (2.2)
2.1 Chau-Keung parametrization
Standard Chau-Keung paramtrization of Pontecorvo-Maki-Nakagawa-Sakata (PMNS) matrix $V_{\text{PMNS}}$ has three Euler angles $\theta_{12},\theta_{23},\theta_{13}$ and Dirac CP-phase $\delta$:
$\begin{split}&V_{\text{CK}}=\begin{pmatrix}c_{12}c_{13}&c_{13}s_{12}&e^{-i\delta}s_{13}\\ -c_{23}s_{12}-c_{12}e^{i\delta}s_{13}s_{23}&c_{12}c_{23}-e^{i\delta}s_{12}s_{13}s_{23}&c_{13}s_{23}\\ s_{12}s_{23}-c_{12}c_{23}e^{i\delta}s_{13}&-c_{12}s_{23}-c_{23}e^{i\delta}s_{12}s_{13}&c_{13}c_{23}\\ \end{pmatrix}=\\ &=\begin{pmatrix}c_{12}&s_{12}&0\\ -s_{12}&c_{12}&0\\ 0&0&1\\ \end{pmatrix}\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&e^{i\delta}\end{pmatrix}\begin{pmatrix}c_{13}&0&s_{13}\\ 0&1&0\\ -s_{13}&0&c_{13}\\ \end{pmatrix}\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&e^{-i\delta}\end{pmatrix}\begin{pmatrix}1&0&0\\ 0&c_{23}&s_{23}\\ 0&-s_{23}&c_{23}\\ \end{pmatrix}\end{split}$ (2.3)
We use the following notation:
$s_{ij}=\sin(\theta_{ij}),\quad c_{ij}=\cos(\theta_{ij}),\quad s2_{ij}=\sin(2\theta_{ij}),\quad s_{\delta}=\sin(\delta),\quad c_{\delta}=\cos(\delta),$ (2.4)
where $i,j\in\{1,2,3\}$.
The complete form of mass matrix (2.2) can be expressed as a sum of Hermitian matrices:
$\begin{split}\mathbb{M}_{\text{flavor}}&=\cfrac{m_{1}}{2}\cdot\begin{pmatrix}2c_{12}^{2}c_{13}^{2}&\begin{split}-s2_{12}c_{13}c_{23}-\\ -e^{-i\delta}\ s2_{13}s_{23}c_{12}^{2}\end{split}&\begin{split}s2_{12}s_{23}c_{13}-\\ -e^{-i\delta}\ s2_{13}c_{12}^{2}c_{23}\end{split}\\ \begin{split}\\ -s2_{12}c_{13}c_{23}-\\ -e^{i\delta}\ s2_{13}s_{23}c_{12}^{2})\end{split}&\begin{split}\\ 2s_{13}^{2}s_{23}^{2}c_{12}^{2}+2s_{12}^{2}c_{23}^{2}+\\ +c_{\delta}s2_{12}s2_{23}s_{13}\end{split}&\begin{split}\\ s_{13}^{2}s2_{23}c_{12}^{2}-s_{12}^{2}s2_{23}+\\ +s2_{12}s_{13}(2c_{\delta}c_{23}^{2}-e^{i\delta})\end{split}\\ \begin{split}\\ s2_{12}s_{23}c_{13}-\\ -e^{i\delta}\ s2_{13}c_{12}^{2}c_{23}\end{split}&\begin{split}\\ s_{13}^{2}s2_{23}c_{12}^{2}-s_{12}^{2}s2_{23}+\\ +s2_{12}s_{13}(2c_{\delta}c_{23}^{2}-e^{-i\delta})\end{split}&\begin{split}\\ 2s_{13}^{2}c_{12}^{2}c_{23}^{2}+2s_{12}^{2}s_{23}^{2}-\\ -c_{\delta}s2_{12}s2_{23}s_{13}\end{split}\\ \end{pmatrix}+\\ &+\cfrac{m_{2}}{2}\cdot\begin{pmatrix}2s_{12}^{2}c_{13}^{2}&\begin{split}s2_{12}c_{13}c_{23}-\\ -e^{-i\delta}s_{12}^{2}s2_{13}s_{23}\end{split}&\begin{split}-s2_{12}c_{13}s_{23}-\\ -e^{-i\delta}s_{12}^{2}s2_{13}c_{23}\end{split}\\ \begin{split}\\ s2_{12}c_{13}c_{23}-\\ -e^{i\delta}s_{12}^{2}s2_{13}s_{23}\end{split}&\begin{split}\\ 2s_{13}^{2}s_{23}^{2}s_{12}^{2}+2c_{12}^{2}c_{23}^{2}-\\ -c_{\delta}s2_{12}s_{13}s2_{23}\end{split}&\begin{split}\\ s_{12}^{2}s_{13}^{2}s2_{23}-c_{12}^{2}s2_{23}-\\ -s2_{12}s_{13}(2c_{\delta}c_{23}^{2}-e^{i\delta})\end{split}\\ \begin{split}\\ -s2_{12}c_{13}s_{23}-\\ -e^{i\delta}s_{12}^{2}s2_{13}c_{23}\end{split}&\begin{split}\\ s_{12}^{2}s_{13}^{2}s2_{23}-c_{12}^{2}s2_{23}-\\ -s2_{12}s_{13}(2c_{\delta}c_{23}^{2}-e^{-i\delta})\end{split}&\begin{split}\\ 2s_{12}^{2}s_{13}^{2}c_{23}^{2}+2c_{12}^{2}s_{23}^{2}+\\ +c_{\delta}s2_{12}s2_{23}s_{13}\end{split}\end{pmatrix}+\\ &+\cfrac{m_{3}}{2}\cdot\begin{pmatrix}2s_{13}^{2}&e^{-i\delta}s2_{13}s_{23}&e^{-i\delta}s2_{13}c_{23}\\ e^{i\delta}s2_{13}s_{23}&2c_{13}^{2}s_{23}^{2}&c_{13}^{2}s2_{23}\\ e^{i\delta}s2_{13}c_{23}&c_{13}^{2}s2_{23}&2c_{13}^{2}c_{23}^{2}\end{pmatrix}\end{split}$ (2.5)
2.2 Numerical estimates
For numerical estimates, we use the standard Chau-Keung parametrization, in which the angles and the CP-violating phase have been measured experimentally. We have 2 different situation for Inverted and Normal Ordering mass states.
Parameter Normal Ordering Inverted Ordering
$\theta_{12}\,/^{\circ}$ $33.68^{+0.73}_{-0.70}$ $33.68^{+0.73}_{-0.70}$
$\theta_{23}\,/^{\circ}$ $43.3^{+1.0}_{-0.8}$ $47.9^{+0.7}_{-0.9}$
$\theta_{13}\,/^{\circ}$ $8.56^{+0.11}_{-0.11}$ $8.59^{+0.11}_{-0.11}$
$\delta_{CP}\,/^{\circ}$ $212^{+26}_{-41}$ $274^{+22}_{-25}$
$\cfrac{\Delta m^{2}_{21}}{10^{-5}\,\text{eV}^{2}}$ $7.49^{+0.19}_{-0.19}$ $7.49^{+0.19}_{-0.19}$
$\cfrac{\Delta m^{2}_{31}}{10^{-3}\,\text{eV}^{2}}$ $+2.513^{+0.021}_{-0.019}$ $-2.409^{+0.020}_{-0.020}$
$\cfrac{\Delta m^{2}_{32}}{10^{-3}\,\text{eV}^{2}}$ $+2.438^{+0.021}_{-0.019}$ $-2.484^{+0.020}_{-0.020}$
Table 1: NuFIT 6.0 (2024): Oscillation parameters IC24 [10]
It is noteworthy that the sum of the squared mass differences $\Delta m_{12}^{2}+\Delta m_{23}^{2}+\Delta m_{31}^{2}=0$ is analogous to the Mandelstam variables satisfying $s+t+u=\sum\limits_{i}m^{2}_{i}$. Consequently, the mass eigenvalues can be represented graphically 1 as a Mandelstam plane, with the area of the central triangle degenerating to zero.
Figure 1: Mandelstam plane analog
In other words, the mass hierarchy can be described by a single angle $\phi$ and a radius $r$.
One of the mixing angles is small compared to the others. However, how can an increase in this angle affect the values of the mass matrix? What happens to the diagonal elements of this matrix? Will they increase monotonically or not, and when might this behavior change?
Normal hierarchy ($\mathbb{M}_{\text{NH}}\times 10^{-2}\text{ eV}$)
$\mathbb{M}_{\text{NH}}(m_{1}=0)\simeq\begin{pmatrix}0.371&-0.119+0.254i&-0.702+0.269i\\ -0.119-0.254i&2.676&2.153\,-0.032i\\ -0.702-0.269i&2.153\,+0.032i&2.831\\ \end{pmatrix}$ (2.6)
Inverse hierarchy ($\mathbb{M}_{\text{IH}}\times 10^{-2}\text{ eV}$):
$\mathbb{M}_{\text{IH}}(m_{3}=0)\simeq\begin{pmatrix}4.822&-0.015-0.539i&-0.06-0.487i\\ -0.015+0.539i&2.29&-2.413-0.005i\\ -0.06+0.487i&-2.413+0.005i&2.781\\ \end{pmatrix}$ (2.7)
Substituting the values for the angles and masses into the expression for the mass, it was observed that for the normal hierarchy (NH), the values of the diagonal elements of the mass matrix form a monotonically increasing sequence. However, upon increasing the mixing angle $\theta_{13}$ and exceeding $\theta_{13}\simeq 19.8^{\circ}$, the sequence ceases to be monotonically increasing. The same occurs when changing the masses: if the absolute difference between $m_{1}$ and $m_{2}$ is larger than their difference from $m_{3}$, the sequence of diagonal elements of the mass matrix also ceases to be monotonically increasing.
For the inverted hierarchy (IH), the sequence of elements on the main diagonal of the mass matrix was initially neither monotonically increasing nor decreasing. However, upon reaching an angle $\theta_{13}\simeq 36.5^{\circ}$ or larger, the values on the main diagonal formed a monotonically increasing sequence:
(3.186, 3.188, $(3.186,\ 3.188,\ 3.517)\cdot 10^{-2}.$ (2.8)
Substituting known mixing angles and masses in (2.2), we obtain the dependence plots of neutrino mass matrix elements on the lightest neutrino mass for both mass hierarchies (AppendixA).
It was observed that for the inverted hierarchy (IH), the real values $\Re(m_{e\mu})=\Re(m_{\mu e})$ decrease sharply as the mass $m_{3}$ approaches its limit:
$\begin{split}&m_{3}=0\times 10^{-2}\text{ eV}\ \sim\ \Re(m_{e\mu})=-15\cdot 10^{-3},\\ &m_{3}=2\times 10^{-2}\text{ eV}\ \sim\ \Re(m_{e\mu})=-4\cdot 10^{-3},\\ &m_{3}=4\times 10^{-2}\text{ eV}\ \sim\ \Re(m_{e\mu})=-6\cdot 10^{-5}.\end{split}$ (2.9)
Analytical formula of $\Re(m_{e\mu})$ for Chau-Keung parametrization:
$\begin{split}\Re(m_{e\mu})=&-\cfrac{m_{1}}{2}\cdot(s2_{12}c_{13}c_{23}+c_{\delta}s2_{13}s_{23}c_{12}^{2})+\\ &+\cfrac{m_{2}}{2}\cdot(s2_{12}c_{13}c_{23}-c_{\delta}s_{12}^{2}s2_{13}s_{23})+\\ &+\cfrac{m_{3}}{2}\cdot c_{\delta}s2_{13}s_{23}\end{split}.$ (2.10)
Plots for both hierarchy (Fig. 2).
Figure 2: Plots of Inverse (left) and Normal (right) Hierarchy as a function of lightest neutrino mass (IH:$m_{3}$, NH:$m_{1}$).
As the lightest neutrino mass increases, an asymptotic approach to zero is observed for the off-diagonal elements. This behavior originates from the following relations for the normal hierarchy:
$m_{2}=m_{1}\cdot\sqrt{1+\alpha_{2}+\alpha_{2}^{2}},\quad m_{3}=m_{1}\cdot\sqrt{1+\alpha_{3}+\alpha_{3}^{2}},$ (2.11)
where $\alpha_{2}=\cfrac{m_{2}-m_{1}}{2m_{1}},\ \alpha_{3}=\cfrac{m_{3}-m_{1}}{2m_{1}}$.
For the inverted hierarchy:
$m_{1}=m_{3}\cdot\sqrt{1+\beta_{1}+\beta_{1}^{2}},\quad m_{2}=m_{3}\cdot\sqrt{1+\beta_{2}+\beta_{2}^{2}},$ (2.12)
where $\beta_{1}=\cfrac{m_{1}-m_{3}}{2m_{3}},\ \beta_{2}=\cfrac{m_{2}-m_{3}}{2m_{3}}$. In the large-mass limit for the lightest neutrino ($m_{1}$ in NH, $m_{3}$ in IH), $\alpha_{i}\to 0$ and $\beta_{i}\to 0$. Consequently, according to (2.2), the flavor mass matrix reduces to the identity form for both hierarchies:
$\lim_{\alpha_{i}\to 0}\mathbb{M}^{\text{NH}}_{\text{flavor}}=m_{1}\cdot\mathbb{I},\quad\lim_{\beta_{i}\to 0}\mathbb{M}^{\text{IH}}_{\text{flavor}}=m_{3}\cdot\mathbb{I}.$ (2.13)
All non-diagonal elements of neutrino mass matrix exhibit power-law decay towards zero without sign inversion. However, $\Re(m_{e\mu})$ displays distinct behavior in both mass hierarchies. For the Normal Hierarchy (NH), we observe an initial local minimum followed by asymptotic convergence to zero. In the Inverted Hierarchy (IH), $\Re(m_{e\mu})$ undergoes sign reversal, reaches a global maximum beyond cosmological bounds on $m_{3}$, and subsequently approaches zero from above. This pronounced peak, while cosmologically unphysical, presents a striking feature. The inverse hierarchy is related to some kind of fine-tunning for matrix. This may be an additional argument in choosing the normal hierarchy (NH).
3 Weak Interaction Process
In the Standard Model, weak interactions involve only left-handed neutrinos. The interaction Lagrangian is:
$\mathcal{L}_{\text{int}}(x)=-\frac{g}{\sqrt{2}}\sum_{\ell}\overline{\ell}(x)\gamma^{\mu}\mathcal{P}_{L}\nu_{\ell}(x)W_{\mu}^{-}(x)-\frac{g}{\sqrt{2}}\sum_{\ell}\overline{\nu}_{\ell}(x)\gamma^{\mu}\mathcal{P}_{L}\ell(x)W_{\mu}^{+}(x),$ (3.1)
where $\mathcal{P}_{L}=(1-\gamma_{5})/2$ are the left-handed projectors, satisfying $\mathcal{P}_{L}\mathcal{P}_{R}=0$. In Feynman diagrams, vertices include $\gamma^{\mu}\mathcal{P}_{L}$. Consequently, in the amplitude squared, terms with constants between vertices vanish due to:
$\mathcal{P}_{L}\cdot C\cdot\gamma^{\nu}\mathcal{P}_{L}=C\cdot\mathcal{P}_{L}\mathcal{P}_{R}\gamma^{\nu}=0.$ (3.2)
3.1 Process $l_{\alpha}^{-}l_{\beta}^{+}\to W^{+}W^{-}$
We consider the process of two leptons interacting to form two W-bosons and evaluate the contribution of oscillations to this local process.
Figure 3: Diagrams of $l_{\alpha}^{-}l_{\beta}^{+}\to W^{+}W^{-}$ for t-channel (3) and u-channel (3).
For the kinematics, we assume negligible lepton masses compared to their energies and negligible neutrino masses. In the center-of-mass frame of the incoming leptons:
$\begin{split}&p_{1}\simeq(|\mathbf{p}_{in}|,\mathbf{p}_{in}),\\ &p_{2}\simeq(|\mathbf{p}_{in}|,-\mathbf{p}_{in}),\\ &p_{3}\simeq p_{4}\simeq(M_{W},0).\end{split}$ (3.3)
Given the propagator form (1.7), the property (3.2), and the kinematics (3.3), the amplitude squared $\mathcal{A}^{2}_{\alpha\beta}$ in the flavor basis simplifies to:
$\mathcal{A}^{2}\simeq\frac{8g^{4}}{M_{W}^{4}}\begin{pmatrix}M_{W}^{4}&\text{tr}^{2}(\mathbb{M})\cdot m_{e\mu}^{2}\\ \text{tr}^{2}(\mathbb{M})\cdot m_{e\mu}^{2}&M_{W}^{4}\end{pmatrix},$ (3.4)
where diagonal elements correspond to the exchange of neutrinos of the same flavor, and off-diagonal elements to oscillations.
The differential cross-section for this process is:
$\frac{d\sigma}{d\Omega}=\frac{\mathcal{A}^{2}}{64\pi^{2}s}\simeq\frac{g^{4}}{16\pi^{2}}\frac{1}{M_{W}^{6}}\begin{pmatrix}M_{W}^{4}&\text{tr}^{2}(\mathbb{M})\cdot m_{e\mu}^{2}\\ \text{tr}^{2}(\mathbb{M})\cdot m_{e\mu}^{2}&M_{W}^{4}\end{pmatrix}.$ (3.5)
The squared amplitude (3.4) reveals the following:
  • Diagonal terms ($\alpha=\beta$): Suppressed by $M_{W}^{2}$, yielding cross-sections $\propto 10^{-26}$ b/sr.
  • Off-diagonal terms ($\alpha\neq\beta$): Suppressed by $\text{tr}^{2}(\mathbb{M})m_{e\mu}^{2}/M_{W}^{6}\sim 10^{-79}$ b/sr due to tiny neutrino masses.
$\frac{d\sigma_{\mu^{+}e^{-}\to W^{+}W^{-}}}{d\sigma_{e^{+}e^{-}\to W^{+}W^{-}}}\simeq\cfrac{\text{tr}^{2}(\mathbb{M})m_{e\mu}^{2}}{M_{W}^{4}}\simeq 3\cdot 10^{-53}$ (3.6)
The obtained cross section is 53 orders of magnitude lower than that of the reference channel while being flavor-conserving, which precludes the possibility of observing such processes in current and future experiments. This result thus validates neglecting the contributions from flavor-changing neutrino exchange interactions in the analysis of $W$-boson production processes in lepton collisions.
3.2 Neutrino Oscillation
Neutrino oscillations significantly affect processes over large space intervals $\mathbf{x}-\mathbf{y}\gg 0$ and positive time difference $x_{0}-y_{0}>0$.
Stray Forward to by deduce calculations witch will be reported elsewhere, considering processes with two flavor changes allows us to reproduce the standard formula for the probability of neutrino oscillations:
$P_{e\to\mu}\propto\cfrac{\sin^{2}(2\theta)}{\text{L}^{2}}\ \sin^{2}\left(\cfrac{\Delta m^{2}}{4E_{\text{avg}}}\text{L}\right),$ (3.7)
where values
$\Delta m^{2}=\text{tr}(\mathbb{M})\sqrt{\text{tr}^{2}(\mathbb{M})-4\det(\mathbb{M})}\quad\text{and}\quad\sin^{2}(2\theta)=\cfrac{2m_{e\mu}}{\sqrt{\text{tr}^{2}(\mathbb{M})-4\det(\mathbb{M})}}$ (3.8)
are expressed through the invariants of the mass matrix. The quantities $L$ are the distance between the source and the detector, and $E_{\text{avg}}$ are the average energies in the detector and the source.
Conclusions
Consideration of processes with violation of lepton number showed that it does not matter whether we work in the mass or flavor states basis. We obtain the same expressions, while in weak interactions flavor states appear.
In this work, we derived the matrix propagator both in its general form and with explicit angular dependence. Also we write explicit mass matrix for 3 flavor through masses and angles. It appears to be qualitatively different in the cases of direct and inverse hierarchy, and the latter seems to exhibit a sort of fine tuning, which may be considered as an extra support in favor of direct hierarchy.
For the local diagram of the flavor-violating process $l_{\alpha}^{-}l_{\beta}^{+}\to W^{+}W^{-}$, we obtained a cross section of $\mathcal{O}(10^{-53})$, indicating no observable contribution to this process.
Coming back to the question posed in the title we may say that there seems to be no definite answer as soon as weak interactions are considered. The appearance of massive and flavor states may be considered as a kind of Bohr’s complementarity. At the same time, consideration of neutrino helicity flip due to interactions with gravity (see e.g. [12, 13, 14]) may require the appearance of definite massive states.
Acknowledgments
We are most indebted to V.A. Naumov for useful discussions and valuable comments.
Appendix A Plots of mass matrix element dependence for NH and IH.
Plots of mass matrix element dependence on $m_{1}$ for Normal Hierarchy
Figure 4: Diagonal elements $m_{ee}$, $m_{\mu\mu}$, $m_{\tau\tau}$ of the neutrino mass matrix in NH versus $m_{1}$
Figure 5: Real and imaginary components of $m_{e\mu}$ in NH as a function of $m_{1}$
Figure 6: Real and imaginary components of $m_{e\tau}$ in NH versus $m_{1}$
Figure 7: Real and imaginary components of $m_{\mu\tau}$ in NH as a function of $m_{1}$
Plots of mass matrix element dependence on $m_{3}$ for Inverted Hierarchy
Figure 8: Diagonal elements $m_{ee}$, $m_{\mu\mu}$, $m_{\tau\tau}$ of the neutrino mass matrix in IH versus $m_{3}$
Figure 9: Real and imaginary components of $m_{e\mu}$ in IH as a function of $m_{3}$
Figure 10: Real and imaginary components of $m_{e\tau}$ in IH versus $m_{3}$
Figure 11: Real and imaginary components of $m_{\mu\tau}$ in IH as a function of $m_{3}$